# Concurrent Balanced Augmented Trees

### v1

@@ Abstract.
Augmentation makes search trees tremendously more versatile, allowing them to support efficient aggregation queries, order-statistic queries, and range queries in addition to insertion, deletion, and lookup. In this paper, we present the first lock-free augmented balanced search tree. Our algorithmic ideas build upon a recent augmented unbalanced search tree presented by Fatourou and Ruppert [DISC, 2024]. We implement both data structures, solving some memory reclamation challenges in the process, and provide an experimental performance analysis of them. We also present optimized versions of our balanced tree that use delegation to achieve better scalability and performance (by more than 2x in some workloads). Our experiments show that our augmented balanced tree is 2.2 to 30 times faster than the unbalanced augmented tree, and up to several orders of magnitude faster than unaugmented trees on 120 threads.
conference: To appear in PPoPP; 2026; Sydney
@@ 1. Introduction
Sets and Key-Value Stores—which support insertions, deletions, and queries—are among the most fundamental and widely used data objects. Sequentially, two efficient classical data structures are used for these objects: Hash Tables and Balanced Search Trees. Hash Tables are fast; but balanced trees are significantly more versatile, since they preserve ordering of keys and support augmentation of the nodes with extra information, to enable many more operations efficiently, including aggregation queries, such as size, sum of values, maximum, minimum, and average; order statistics, such as finding the $i$th smallest (or largest) key in the set, or the rank of a given key; and range queries that list or aggregate keys in a given range. While augmentation of sequential balanced trees is an indispensable and widely-used technique that is discussed in standard undergraduate algorithms textbooks (Cormen et al., 2022; Goodrich and Tamassia, 2014; Sanders et al., 2019; Sedgewick and Wayne, 2011), the technique has evaded concurrent implementation until now. We design and implement the first lock-free augmented balanced search tree and demonstrate the efficiency of our data structure empirically.
@@ 1.1. Approach and Challenges
Most BST operations take time proportional to the tree’s height, which can be linear in the number of keys in the tree. Hence, balanced BSTs, which guarantee the height is logarithmic in the number of keys, are often preferable. We show how to extend Fatourou and Ruppert’s augmentation technique to get a lock-free augmented balanced BST. This requires coping with rotations (rebalancing operations), which can change the structure of the tree at any location, whereas the original paper dealt only with insertions and deletions of leaves. We apply our extension of the augmentation technique to Brown, Ellen and Ruppert’s lock-free implementation (Brown et al., 2014) of a chromatic BST  (Nurmi and Soisalon-Soininen, 1996), which provides balancing guarantees. We call our data structure BAT (Balanced Augmented Tree).
We also show how to add memory reclamation to both the augmented unbalanced BST (Fatourou and Ruppert, 2024) and the BAT. This required solving several novel challenges, since some shared objects can be reached via multiple paths, so it requires care to track when they have been fully removed from the data structure. We present a lightweight method for this tracking that avoids the overhead of reference counters. A crucial observation is that some objects are safe to free even while still reachable from the root if we can guarantee that no operation will access them.
@@ 1.2. Our Contributions
  • Lock-Free BAT. We design the first lock-free balanced augmented search tree data structure.
@@
  • Performance. Our experiments show that our BAT is between 2.2 and 30 times faster than the augmented unbalanced tree (Fatourou and Ruppert, 2024) across all our experiments. Compared to the fastest unaugmented concurrent tree (Blelloch and Wei, 2024), BAT is up to several orders of magnitude faster in workloads with order-statistic queries or large range queries.
@@ 2. Related Work
Three papers on augmenting concurrent (unbalanced) search trees appeared in 2024: Fatourou and Ruppert (FR) (Fatourou and Ruppert, 2024), Kokorin, Yudov, Aksenov and Alistarh (KYAA) (Kokorin et al., 2024), and Sela and Petrank (SP) (Sela and Petrank, 2024). Our approach extends that of FR (detailed in Section 3.2), which is the most general of the three and also the only one that results in a lock-free data structure. Our safe memory reclamation technique from Section 6 can be applied to their approach too.
KYAA use a lock-based approach in which each node has an associated FIFO queue, and before accessing a node, an operation must join the corresponding queue and help all operations ahead of it before reading or writing to it. KYAA’s approach is specifically designed for order-statistic trees, and it is not clear how to generalize it to other augmentations.
SP also gave a lock-based augmented tree that supports aggregating functions formed using Abelian group operators (i.e., a generalization of augmenting nodes with the sizes of their subtrees). In their approach, update operations announce themselves with timestamps, and each query must gather information from ongoing updates with smaller timestamps than the query using the multiversioning approach of Wei et al. (Wei et al., 2021).
Our augmentation scheme, like FR’s, has the bonus property of providing simple atomic snapshots of the set of keys in the BST. Taking snapshots of shared data structures has received much attention recently (Fatourou et al., 2019; Nelson-Slivon et al., 2022; Prokopec et al., 2012; Wei et al., 2021; Jayanti et al., 2024; Bashari et al., 2024; Jayanti and Jayanti, 2025a, b). Naïve but inefficient algorithms for order-statistic queries can use such snapshots. For example, one can count the keys in a given range by taking a snapshot and traversing all keys in the range. This takes time linear in the number of keys in the range. Our augmented trees can answer these queries much more efficiently by traversing just two paths of the BST in time proportional to the tree’s height. In our experimental analysis we compare the performance of these two approaches to answering such queries, where the snapshots are provided by the general technique of Wei et al. (Wei et al., 2021).
Our BAT data structure builds on Nurmi and Soisalon-Soininen’s chromatic tree (Nurmi and Soisalon-Soininen, 1996), which was implemented in a lock-free manner by Brown, Ellen and Ruppert (Brown et al., 2014). The latter implementation is described in Section 3.1.
@@ 3.1. Lock-free Chromatic Trees
For example, Figure 1 depicts one of the 22 possible rotations, called RB1; the nodes of the old patch, marked with $\times$’s, are removed from the tree and the new patch consists of two new nodes, marked with ++’s. The rotation does not change the subtrees labelled by Greek letters (some may be empty subtrees). To perform this rotation, a process first does an LLX on and reads the shaded nodes $A,B,C$. (If any node of these nodes is Finalized, the rotation is aborted, since some other concurrent process has updated this portion of the tree.) The process creates new nodes $B^{\prime}$ and $C^{\prime}$, with the same keys as $B$ and $C$, using information returned by the LLX operations on $B$ and $C$. Finally, the process uses SCX to atomically update the child pointer of $A$ to point to $C^{\prime}$ and simultaneously finalize nodes $B$ and $C$, which have been removed from the tree. If this SCX succeeds, no other updates have modified $A$, $B$ or $C$ after the LLX operations on them, ensuring the modification makes the atomic change shown in Figure 1. Figure 2 shows other examples of tree updates.
@@ 3.2. Augmenting Search Trees
Each node of the search tree has a pointer to a version object, which stores a version of the supplementary fields of the node. The version of a node $x$ is updated by performing a refresh on $x$. The refresh reads $x$’s children’s versions $\mbox{{v}}_{\ell}$ and $\mbox{{v}}_{r}$, computes the value of $x$’s supplementary fields, creates a new version object $\mbox{{v}}^{\prime}$ constructed using this information and finally performs a CAS to swing $x$’s version pointer to $\mbox{{v}}^{\prime}$. In addition to the supplementary fields, $\mbox{{v}}^{\prime}$ stores the key of $x$ and pointers to $\mbox{{v}}_{\ell}$ and $\mbox{{v}}_{r}$. Thus, the version objects themselves form a BST (called the version tree) that mirrors the structure of the original tree (called the node tree). See Figure 4.
@@
30:  Version* $v\leftarrow\mbox{{Root}}.\mbox{{version}}$ $\triangleright$ Start at the root
@@
32:    $v\leftarrow(\mbox{{key}}<v.\mbox{{key}}\ ?\ v.\mbox{{left}}:v.\mbox{{right}})$
@@
34:  return ($v.\mbox{{key}}=\mbox{{key}}$)
@@ 4. Lock-Free Balanced Augmented Tree
Recall that all changes to the chromatic tree of (Brown et al., 2014) are performed by an SCX that replaces one patch of the BST with a new patch consisting of new nodes and simultaneously finalizes the nodes of the replaced patch. Whenever a new patch is created for an insertion, deletion or rebalancing step, it uses the following rules to initialize its nodes’ versions. (We use size as an example augmentation; supplementary fields required by any other augmentation could be used instead.)
@@
To facilitate this, Propagate uses a thread-local stack to keep track of the nodes that it should refresh. The operation first pushes the internal nodes that were visited to get from the root to the leaf $\ell$. If there were no concurrent modifications to the node tree, Propagate could simply retrace its steps, refreshing each node it pops off the stack. However, concurrent operations on the node tree may have added nodes or replaced nodes that appear in this stack. Since each refresh transmits information only from children to parent, any gap in the chain of refreshed nodes would prevent information about the update from reaching the root. Before proceeding to the top node $x$ on the stack, Propagate checks whether it has already refreshed $x$’s current child. If so, it pops $x$ and refreshes it. Otherwise, it traverses down the tree from $x$ (in the direction towards the update’s key), pushing nodes onto the stack until it pushes a node $y$ whose child has been refreshed (or is a leaf). It then pops $y$ and refreshes it. Propagate repeats this process until the root is refreshed.
@@
All newly created internal nodes are initialized with nil version pointers (Definition 1) to indicate that their supplementary fields have not yet been computed. If an update op ever reads a node $x$’s nil version pointer, op performs a refresh to fix $x.\mbox{{version}}$ using information read from $x$’s children’s versions. However, the version pointers of $x$’s children may themselves be nil. In this case, op recursively tries to fix the version pointers of the children by reading $x$’s grandchildren. This process goes down the tree recursively until it finds nodes whose version pointers are non-nil. (This is guaranteed to happen, since leaf nodes never have nil version pointers.) Then, the recursion stops and the version pointers of all nodes visited during the recursion are set to non-nil values.
@@
Description of the Pseudocode. Figure 3 presents the pseudocode for BAT. It uses two types of objects. A Node object represents a tree node and stores a key key, a weight (used for rebalancing), the pointers left and right to the Node’s children, and the version pointer. The first four fields of a node constitute an LLX/SCX record. A Version object represents one version of the supplementary fields of a node. It stores a key key, the size of the subtree rooted at the version, and the pointers to the left and right children in the version tree.
BAT’s Insert and Delete operations simply call CTInsert and CTDelete to perform the update on the node tree, with one change: the version field of any newly-allocated node is initialized using the version initialization rules of Definition 1 (line 18 or line 25). Then, the operation invokes Propagate (line 19 or line 26). Even unsuccessful updates must call Propagate. For example, a process $p$’s Delete(key) might fail because key was deleted from the node tree by a concurrent process $q$ that has not propagated its deletion to the root, so $p$ must ensure $q$’s deletion is propagated before $p$ returns false.
The loop at lines 3951 of Propagate ensures that information about an update operation reaches the root by performing a double Refresh (lines 4748) on a sequence of nodes, as described above. It uses the thread-local stack to determine which nodes to refresh. Initially, stack is empty. In the first iteration of the outer loop, the inner loop at lines 4144 pushes all nodes visited when searching for key key onto stack. Since CTInsert and CTDelete also execute a search for key, we could have those routines store the nodes they visit in stack before calling Propagate as an optimization for performance.
@@
A Refresh on node $x$ reads $x$’s left and right child pointers and the children’s version pointers (lines 5657 and 6465). If either child’s version is nil (lines 58 and 66), it recursively refreshes that child (lines 59 and 67). When all the recursive calls return, Refresh has all the needed information to refresh $x$. Then, it allocates a new version object initializing its fields appropriately (line 72), and attempts to change the version pointer of $x$ to point to this new Version (line 73).
@@ 5. Reducing Contention via Delegation
In our first implementation, called BAT-Del, an instance $P$ of Propagate delegates its work after failing both of its attempts to refresh a node $x$ at lines 47 and 48 of Figure 3. This means some other successful Refresh on $x$ occurs between the start of $P$’s first failed Refresh and the end of $P$’s second failed Refresh. Let $R_{\ell}$ be the last such successful Refresh and $P_{\ell}$ be the Propagate that called $R_{\ell}$. After $P$’s second failed Refresh, if $x$ is not finalized (i.e., it is still in the tree), $P$ delegates the rest of its work to $P_{\ell}$ by waiting for $P_{\ell}$ to complete before returning. When $P_{\ell}$ completes, all of operations $P$ was attempting to propagate will have reached the root.
@@
We next describe how $P$ synchronizes with $P_{\ell}$. Each call to Propagate creates a PropStatus object, which stores a boolean value done indicating whether or not the Propagate has finished, and a pointer to another PropStatus if the Propagate has delegated its work (if not, this pointer is nil). Each version stores a pointer to the PropStatus of the Propagate that created it. When a Propagate $P$ fails the CAS of its second Refresh on a node, the CAS returns the version written by the last successful Refresh $R_{\ell}$, and $P$ can delegate to that version’s PropStatus, which belongs to $_{\ell}. PP$ waits for $P_{\ell}$ to finish by spinning on the done field in the PropStatus object. There may be a chain of delegations, so to avoid waiting for the done flag to propagate down the chain, $P$ can find the head of the chain using the pointers in the PropStatus objects, and directly wait on the head of the chain.
@@
For correctness, we distinguish between top-level Refreshes called by Propagate and recursive Refreshes called by Refresh to fix nil version pointers. We need to ensure that a top-level Refresh cannot fail (and thus delegate) due to a recursive Refresh. We do this by making the CAS in recursive Refreshes only change version pointers from nil to non-nil and the CAS in top-level Refreshes change version pointers only from non-nil to non-nil. We accomplish this by creating two separate refresh functions that only differ in their first few steps. A recursive Refresh begins by reading the node’s version pointer and returning if it is non-nil. A top-level Refresh begins by reading the node’s version pointer and, if it is nil, calling a recursive refresh and rereading the node’s version pointer (which is now guaranteed to no longer be nil). The remaining steps for both versions of Refresh are the same as in Figure 3, with all calls to Refresh going to the recursive version. Delegating due to a recursive Refresh is dangerous because a Propagate may perform recursive Refreshes on nodes outside of its search path. For example, when a new patch is installed, a Propagate might recursively Refresh every node in the new patch, even though many are on different search paths.
@@
In general, it is safe for a Propagate $P_{1}$ to delegate to another Propagate $P_{2}$ at node $x$ if (1) $x$ is still reachable from the root, (2) $x$ is on both their search paths, and (3) the Refresh of $P_{2}$ saw all the arrival points $A$ that $P_{1}$ was attempting to propagate to $x$. Properties (1) and (2) are important for arguing that $P_{2}$ will perform a sequence of Refreshes from $x$ (or from a new node replacing $x$, which inherits all the arrival points in $A$) to the root. This sequence of refreshes has the same effect as continuing $P_{1}$ because property (3) ensures they will bring the arrival points in $A$ to the root.
@@
In our experiments, BAT-Del improves performance by more than a factor of two in update-heavy workloads. We develop a more optimized version called BAT-EagerDel, which further increases the frequency of delegation. BAT-EagerDel uses the same delegation mechanism, but does so after just one failed Refresh. To make this safe, we had to modify a successful top-level Refresh on a node $x$ to re-read the version pointers in $x$’s children and make sure they have not changed after they were last read on line 57 or  60, and line  65 or  68 of the Refresh. If they have changed, the successful Refresh repeats from the beginning (line 54) until either it fails a CAS and sees that $x$ has not been removed (in which case it delegates) or it succeeds in a CAS and sees that the version pointers in $x$’s children have not changed.
@@
Both delegation techniques described above are blocking: if a thread that has undertaken the work of other threads stalls, it prevents the other threads from completing. We can make both BAT-Del and BAT-EagerDel non-blocking by adding a timeout, after which the waiting process resumes its propagation to the root itself. Our implementations in the experiments section include this timeout, tuned for the common case where there are no stalled threads.

### v2

@@ Abstract.
Augmentation makes search trees tremendously more versatile, allowing them to support efficient aggregation queries, order-statistic queries, and range queries in addition to insertion, deletion, and lookup. In this paper, we present the first lock-free augmented balanced search tree supporting generic augmentation functions. Our algorithmic ideas build upon a recent augmented unbalanced search tree presented by Fatourou and Ruppert [DISC, 2024]. We implement both data structures, solving some memory reclamation challenges in the process, and provide an experimental performance analysis of them. We also present optimized versions of our balanced tree that use delegation to achieve better scalability and performance (by more than 2x in most workloads). Our experiments show that our augmented balanced tree completes updates 2.2 to 30 times faster than the unbalanced augmented tree, and outperforms unaugmented trees by up to several orders of magnitude on 120 threads.
Lock-free, Balanced BST, Augmentation, Ordered Statistics Queries, Range Query, Delegation
copyright: ccdoi: 10.1145/3774934.3786437journalyear: 2026isbn: 979-8-4007-2310-0/2026/01conference: Proceedings of the 31st ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming; January 31 – February 4, 2026; Sydney, NSW, Australiabooktitle: Proceedings of the 31st ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming (PPoPP ’26), January 31 – February 4, 2026, Sydney, NSW, Australiaccs: Computing methodologies Concurrent algorithms
@@ 1. Introduction
Sets and Key-Value Stores—which support insertions, deletions, and queries—are among the most fundamental and widely used data objects. Sequentially, two efficient classical data structures are used for these objects: Hash Tables and Balanced Search Trees. Hash Tables are fast; but balanced trees are significantly more versatile, since they preserve ordering of keys and support augmentation of the nodes with extra information, to enable many more operations efficiently, including aggregation queries, such as size, sum of values, maximum, minimum, and average; order statistics, such as finding the $i$th smallest (or largest) key in the set, or the rank of a given key; and range queries that list or aggregate keys in a given range. While augmentation of sequential balanced trees is an indispensable and widely-used technique that is discussed in standard undergraduate algorithms textbooks (Cormen et al., 2022; Goodrich and Tamassia, 2014; Sanders et al., 2019; Sedgewick and Wayne, 2011), the technique has evaded concurrent implementation until now. We provide the first lock-free augmented balanced search tree with unrestricted augmentation and demonstrate its efficiency empirically.
@@ 1.1. Approach and Challenges
Most BST operations take time proportional to the tree’s height, which can be linear in the number of keys in the tree. Hence, balanced BSTs, which guarantee the height is logarithmic in the number of keys, are often preferable. We show how to extend Fatourou and Ruppert’s augmentation technique to get a lock-free augmented balanced BST. This requires coping with rotations (rebalancing operations), which can change the structure of the tree at any location, whereas the original paper dealt only with insertions and deletions of leaves. We apply our extension of the augmentation technique to Brown, Ellen and Ruppert’s lock-free implementation (Brown et al., 2014) of a chromatic BST  (Nurmi and Soisalon-Soininen, 1996), which provides balancing guarantees. We call our data structure BAT (Balanced Augmented Tree). We also show how to add memory reclamation to both the augmented unbalanced BST (Fatourou and Ruppert, 2024) and the BAT.
@@ 1.2. Our Contributions
  • Lock-Free BAT. The first lock-free balanced augmented search tree supporting generic augmentation function.
@@
  • Performance. Our experiments show that our BAT performs updates between 2.2 and 30 times faster than the augmented unbalanced tree (Fatourou and Ruppert, 2024) while maintaining slightly better range query performance across all our experiments. Compared to the fastest unaugmented concurrent tree (Blelloch and Wei, 2024), BAT is up to several orders of magnitude faster in workloads with order-statistic queries or large range queries.
@@ 2. Related Work
Three papers on augmenting concurrent search trees appeared in 2024: Fatourou and Ruppert (FR) (Fatourou and Ruppert, 2024), Sela and Petrank (SP) (Sela and Petrank, 2024), and Kokorin, Yudov, Aksenov and Alistarh (KYAA) (Kokorin et al., 2024). FR and SP focus on unbalanced concurrent search trees, whereas KYAA supports balanced search trees. Our approach extends that of FR (detailed in Section 3.2), which is lock-free and the most general of the three in terms of the augmentation functions it supports.
SP gave a lock-based augmented tree that supports aggregating functions formed using abelian group operators (i.e., a generalization of augmenting nodes with the sizes of their subtrees). In contrast, our approach imposes no such restriction on the augmentation. In their approach, update operations announce themselves with timestamps, and each query must gather information from ongoing updates with smaller timestamps than the query using multiversioning.
KYAA use a wait-free hand-over-hand helping approach in which each node has an associated FIFO queue. Before accessing a node, an operation must join the corresponding queue and help all operations ahead of it before reading or writing to the node. Their approach applies to balanced trees and generalizes to handle a similar class of aggregation functions as SP’s approach.
Our augmentation scheme, like FR’s, has the bonus property of providing simple atomic snapshots of the set of keys in the BST. Taking snapshots of shared data structures has received much attention recently (Fatourou et al., 2019; Nelson-Slivon et al., 2022; Prokopec et al., 2012; Wei et al., 2021; Jayanti et al., 2024; Bashari et al., 2024; Jayanti and Jayanti, 2025a, b; Blelloch and Wei, 2024). Naïve but inefficient algorithms for order-statistic queries can use such snapshots. For example, one can count the keys in a given range by taking a snapshot and traversing all keys in the range. This takes time linear in the number of keys in the range. Our augmented trees can answer these queries much more efficiently by traversing just two paths of the BST in time proportional to the tree’s height. In our experimental analysis we compare the performance of these two approaches to answering such queries, where the snapshots are provided by the general technique of Wei et al. (Wei et al., 2021). Our BAT data structure builds on Nurmi and Soisalon-Soininen’s chromatic tree (Nurmi and Soisalon-Soininen, 1996), which was implemented in a lock-free manner by Brown, Ellen and Ruppert (Brown et al., 2014).
Manohar, Wei and Blelloch (Manohar et al., 2025) presented two techniques for designing lock-free augmented trees that support low-dimensional k-nearest neighbor queries. Their techniques are optimized for the shallow-effect augmentation setting where each update operation typically only changes a few augmented values near the updated leaf rather than changing all of the augmented values up to the root.
@@ 3.1. Lock-free Chromatic Trees
For example, Figure 1 depicts one of the 22 possible rotations in the chromatic tree from (Brown et al., 2014), called RB1; the nodes of the old patch, marked with $\times$’s, are removed from the tree and the new patch consists of two new nodes, marked with ++’s. The rotation does not change the subtrees labelled by Greek letters (some may be empty subtrees). To perform this rotation, a process first does an LLX on and reads the shaded nodes $A,B,C$. (If any node of these nodes is Finalized, the rotation is aborted, since some other concurrent process has updated this portion of the tree.) The process creates new nodes $B^{\prime}$ and $C^{\prime}$, with the same keys as $B$ and $C$, using information returned by the LLX operations on $B$ and $C$. Finally, the process uses SCX to atomically update the child pointer of $A$ to point to $C^{\prime}$ and simultaneously finalize nodes $B$ and $C$, which have been removed from the tree. If this SCX succeeds, no other updates have modified $A$, $B$ or $C$ after the LLX operations on them, ensuring the modification makes the atomic change shown in Figure 1. Figure 2 shows other examples of tree updates.
@@ 3.2. Augmenting Search Trees
Each node of the BST has a pointer to a version object, which stores a version of the node’s supplementary fields. A node $x$’s version is updated by performing a refresh on $x$. The refresh reads $x$’s children’s versions $\mbox{{v}}_{\ell}$ and $\mbox{{v}}_{r}$, computes the value of $x$’s supplementary fields, creates a new version object $\mbox{{v}}^{\prime}$ constructed using this information and finally performs a CAS to swing $x$’s version pointer to $\mbox{{v}}^{\prime}$. In addition to the supplementary fields, $\mbox{{v}}^{\prime}$ stores the key of $x$ and pointers to $\mbox{{v}}_{\ell}$ and $\mbox{{v}}_{r}$. Thus, the version objects themselves form a BST (called the version tree) that mirrors the structure of the original tree (called the node tree). See Figure 4.
@@
30:  Version* $\mbox{{v}}\leftarrow\mbox{{Root}}.\mbox{{version}}$ $\triangleright$ Start at the root
@@
32:    $\mbox{{v}}\leftarrow(\mbox{{key}}<\mbox{{v}}.\mbox{{key}}\ ?\ \mbox{{v}}.\mbox{{left}}:\mbox{{v}}.\mbox{{right}})$
@@
34:  return ($\mbox{{v}}.\mbox{{key}}=\mbox{{key}}$)
@@ 4. Lock-Free Balanced Augmented Tree
All changes to the chromatic tree of (Brown et al., 2014) are performed by an SCX that replaces one patch of the BST with a new patch consisting of new nodes and simultaneously finalizes the nodes of the replaced patch. Whenever a new patch is created for an insertion, deletion or rebalancing step, it uses the following rules to initialize its nodes’ versions. (We use size as an example augmentation; supplementary fields required by any other augmentation could be used instead.)
@@
To facilitate this, Propagate uses a thread-local stack to keep track of the nodes that it should refresh. The operation first pushes the internal nodes that were visited as it traverses from the root to the leaf $\ell$. If there were no concurrent modifications to the node tree, Propagate could simply retrace its steps, refreshing each node it pops off the stack. However, concurrent operations on the node tree may have added nodes or replaced nodes that appear in this stack. Since each refresh transmits information only from children to parent, any gap in the chain of refreshed nodes would prevent information about the update from reaching the root. Before proceeding to the top node $x$ on the stack, Propagate checks whether it has already refreshed $x$’s current child. If so, it pops $x$ and refreshes it. Otherwise, it traverses down the tree from $x$ (in the direction towards the update’s key), pushing nodes onto the stack until it pushes a node $y$ whose child has been refreshed (or is a leaf). It then pops $y$ and refreshes it, repeating this process until the root is refreshed.
@@
All newly created internal nodes are initialized with nil version pointers (Definition 1) to indicate that their supplementary fields have not yet been computed. If an update op ever reads a node $x$’s nil version pointer, op performs a refresh to fix $x.\mbox{{version}}$ using information read from $x$’s children’s versions. If the version pointers of $x$’s children are themselves nil, op recursively tries to fix the children’s version pointers by reading $x$’s grandchildren. This process goes down the tree recursively until it finds nodes whose version pointers are non-nil. (This is guaranteed to happen, since leaf nodes never have nil version pointers.) Then, the recursion stops and the version pointers of all nodes visited during the recursion are set to non-nil values.
@@
Description of the Pseudocode. Pseudocode for BAT is in Figure 3. It uses two types of objects. A Node object represents a tree node and stores a key key, a weight (used for rebalancing), the pointers left and right to the Node’s children, and the version pointer. The first four fields of a node constitute an LLX/SCX record. A Version object represents one version of the supplementary fields of a node. It stores a key key, the size of the subtree rooted at the version, and the pointers to the left and right children in the version tree.
BAT’s Insert and Delete operations simply call CTInsert and CTDelete to perform the update on the node tree, with one change: the version field of any newly-allocated node is initialized using rules of Definition 1 (line 18 or line 25). Then, the operation invokes Propagate (line 19 or line 26). Even unsuccessful updates must call Propagate. For example, a process $p$’s Delete(key) might fail because key was deleted from the node tree by a concurrent process $q$ that has not propagated its deletion to the root, so $p$ must ensure $q$’s deletion is propagated before $p$ returns false.
The loop at lines 3951 of Propagate ensures that information about an update operation reaches the root by performing a double Refresh (lines 4748) on a sequence of nodes, as described above. It uses the thread-local stack to determine which nodes to refresh. Initially, stack is empty. In the first iteration of the outer loop, the inner loop at lines 4144 pushes all nodes visited when searching for key key onto stack. Since CTInsert and CTDelete also execute a search for key, we could have those routines store the nodes they visit in stack before calling Propagate as an optimization.
@@
A Refresh on node $x$ reads $x$’s left and right child pointers and the children’s version pointers (lines 5562 and 6370). If either child’s version is nil (lines 58 and 66), it recursively refreshes that child (lines 59 and 67). The loops ensure that the node $x_{\ell}$ or $x_{r}$ was still a child of $x$ when its version field was read; if not, the loop rereads $x$’s child and its version field. When all the recursive calls return, Refresh has all the needed information to refresh $x$. Then, it allocates a new version object initializing its fields appropriately (line 72), and attempts to change the version pointer of $x$ to point to this new Version (line 73).
@@ 5. Reducing Contention via Delegation
In our first implementation, called BAT-Del, an instance $P$ of Propagate delegates its work if failing both of its attempts to refresh a node $x$ at lines 47 and 48 of Figure 3. This means some other successful Refresh on $x$ occurs between the start of $P$’s first failed Refresh and the end of $P$’s second failed Refresh. Let $R_{\ell}$ be the last such successful Refresh and $P_{\ell}$ be the Propagate that called $R_{\ell}$. After $P$’s second failed Refresh, if $x$ is not finalized (i.e., it is still in the tree), $P$ delegates the rest of its work to $P_{\ell}$ by waiting for $P_{\ell}$ to complete before returning. We say that the delegation happens at $R_{\ell}$’s successful CAS. When $P_{\ell}$ completes, all of the operations $P$ was attempting to propagate will have arrived at the root.
@@
We next describe how $P$ synchronizes with $P_{\ell}$. Each call to Propagate creates a PropStatus object, which stores a boolean value done indicating whether or not the Propagate has finished and a pointer to another PropStatus if the Propagate has delegated its work (if not, this pointer is nil). Each version stores a pointer to the PropStatus of the Propagate that created it. When a Propagate $P$ fails the CAS of its second Refresh on a node, the CAS returns the version written by the last successful Refresh $R_{\ell}$, and $P$ can delegate to that version’s PropStatus, which belongs to $_{\ell}. PP$ waits for $P_{\ell}$ to finish by spinning on the done field in the PropStatus object. There may be a chain of delegations, so to avoid waiting for the done flag to propagate down the chain, $P$ can find the head of the chain using the pointers in the PropStatus objects, and directly wait on the head of the chain.
@@
For correctness, we distinguish between top-level Refreshes called by Propagate and recursive Refreshes called by Refresh to fix nil version pointers. We need to ensure that a top-level Refresh cannot fail (and thus delegate) due to a recursive Refresh. We do this by making the CAS in recursive Refreshes only change version pointers from nil to non-nil and the CAS in top-level Refreshes change version pointers only from non-nil to non-nil. We accomplish this by creating two separate refresh functions that only differ in their first few steps. A recursive Refresh begins by reading the node’s version pointer and returning if it is non-nil. A top-level Refresh begins by reading the node’s version pointer and, if it is nil, calling a recursive refresh and rereading the node’s version pointer (which is now guaranteed to no longer be nil). The remaining steps for both versions of Refresh are the same as in Figure 3. Delegating due to a recursive Refresh is dangerous because a Propagate may perform recursive Refreshes on nodes outside of its search path. For example, when a new patch is installed, a Propagate might recursively Refresh every node in the new patch, even though many are on different search paths.
@@
Next, we sketch why the delegation scheme is correct. In general, it is safe for a Propagate $P_{1}$ to delegate to another Propagate $P_{2}$ at node $x$ if at the time of delegation (1) $x$ is reachable from the root, (2) $x$ is on the search paths of both $P_{1}$ and $P_{2}$, (3) the Refresh of $P_{2}$ sees the set $A$ of all arrival points that $P_{1}$ was attempting to propagate to $x$, and (4) none of the nodes above $x$ are in $P_{2}$’s refreshed set. Properties (1) and (2) are important for arguing that $P_{2}$ will perform a sequence of Refreshes from $x$ (or from a new node replacing $x$, which inherits all the arrival points in $A$) to the root. This sequence of refreshes has the same effect as continuing $P_{1}$ because properties (3) and (4) ensure they will bring the arrival points in $A$ to the root.
Together, these properties can be used to prove the following invariant by induction over the execution history. Consider a Propagate $P$ called by an update operation $U$ with key $k$. Let $x_{C}$ be the top-most node (the node closest to the root) on the search path for $k$ that $U$ has arrived at. The invariant is that at every configuration $C$ during $P$, if $U$ has not arrived at the root, then the Propagate $Q$ at the end of $P$’s delegation chain is at or below $x_{C}$, meaning that $x_{C}$ is on $Q$’s search path and no node between the root (inclusive) and $x_{C}$ (exclusive) is in $Q$’s refreshed set. This invariant ensures that $U$ arrives at the root before the root is added to $Q$’s refreshed set. Since $P$ waits for $Q$, and $Q$ only returns when the root is added to $Q$’s refreshed set, $U$ must arrive at the root before $P$ returns. Thus, BAT-Del preserves the correctness of the original algorithm.
@@
In our experiments, BAT-Del improves performance by more than a factor of two in update-heavy workloads. We develop a more optimized version called BAT-EagerDel, which further increases the frequency of delegation. BAT-EagerDel uses the same delegation mechanism, but changes the Propagate function so that it is safe to delegate if just one Refresh $R$ fails. Let $R_{\ell}$ be the Refresh that caused $R$ to fail. When $R$ delegates to $_{\ell}, RR_{\ell}$ might not have seen all the arrival points that $R$ was attempting to propagate to $x$, violating property (3). To restore the property, we modify the Propagate function so that it keeps calling Refresh on the same node until either (1) the Refresh fails and the node is not finalized (in which case it delegates to the Propagate that caused this Refresh to fail), or (2) the Refresh succeeds and the version pointers in the node’s children are the same before and after the Refresh. Suppose a Propagate $P_{1}$ delegates to a Propagate $P_{2}$ at node $x$. Condition (2) ensures that $P_{2}$ continues to call Refresh $x$ until it sees that the version pointers in $x$’s children do not change during the execution of such a Refresh, at which point $P_{2}$ is guaranteed to have seen all of the arrival points $P_{1}$ was attempting to propagate to $x$, thus restoring property (3). The correctness of BAT-EagerDel uses similar arguments as those for BAT-Del.
@@
Both delegation techniques described above are blocking: if a thread that has undertaken the work of other threads stalls, it prevents the other threads from completing. We can make both BAT-Del and BAT-EagerDel non-blocking by adding a timeout, after which the waiting process resumes its propagation to the root itself.